# Distribution of mixed character sums and extremal problems for Littlewood polynomials

### v1

@@ 2.1. Proofs of Theorems 1 and 2
for all (sufficiently large) $k$ and for all $s$ and $t$ and if the finite dimensional distributions of $L_{k}(t)$ tend to some limits, then $L_{k}(t)$ converges weakly to some random process $L$.
Instead of working directly with the random processes $F_{k,q,\alpha,\beta}$ and $G_{q,\alpha,\beta}$ we define truncated versions $\tilde{F}_{k,q,\alpha,\beta}$ and $\tilde{G}_{q,\alpha,\beta}$ and prove in Proposition 4 that a sequence of these truncated random processes converge if and only if the untruncated processes converge. This is a technical convenience, but it also somewhat necessary; the moments of the untruncated random processes will not converge without removal of the trivial character and moreover, the size of the truncation parameter will play a role in our counting arguments handling various ranges of $k_{q}.$
@@
Finally, to apply Proposition 2, we prove that all mixed moments of all of finite dimentional distributions of $\tilde{F}_{k,q,\alpha,\beta}$ tend to those of $F_{k,\alpha,\beta}$ when $k$ is fixed (Proposition 5) or tend to those of $F_{\alpha,\beta}$ when $k$ tends to infinity, after possibly removing some primes (Proposition 6), in each of the cases (2), (3), and (4) of Theorem 1. These propositions also include bounds for the moments which show that the distributions are determined by the moments (Carleman condition), yielding a proof of Theorem 1. Similarly, we deal with the quadratic case in Proposition 7, completing the proof of Theorem 2.
@@ 3. The random processes and truncation
$\tilde{f}(n)=\frac{1}{2}\lim_{\varepsilon\rightarrow 0}\big(f(n+\varepsilon)+f(n-\varepsilon)\big).$
@@
Equipped with these lemmas we are in a position to state and prove our general formula for incomplete shifted character sums.
@@
For a primitive Dirichlet character $\chi$ mod $q$ and for $\alpha,\beta\in\mathbb{R}$ we have
@@
In particular, for any $\alpha$ and $\beta$ such that $\alpha q$ and $(\alpha+\beta)q$ are not integers,
@@
$\sum_{\alpha q\leq n\leq\alpha q+\beta q}\chi(n)e(n\theta)=e(\alpha q\theta)\frac{\tau(\chi)}{2\pi i}\sum_{m\in\mathbb{Z}}\frac{e(-\alpha m)(e(\beta(\theta q-m))-1)}{\theta q-m}\overline{\chi}(m).$
Now write $\theta=(k+t)/q$, with $k=\left\lfloor q\theta\right\rfloor$ and $t=\left\{\theta\right\}$. Then
$\begin{split}e(-\alpha q\theta)\sum_{\alpha q\leq n\leq\alpha q+\beta q}\chi(n)e(n\theta)=&\frac{\tau(\chi)}{2\pi i}\sum_{m\in\mathbb{Z}}\frac{e(-\alpha m)(e(\beta(k+t-m))-1)}{k+t-m}\overline{\chi}(m)\\ =&\frac{\tau(\chi)}{2\pi i}\sum_{l\in\mathbb{Z}}\frac{e(-\alpha(k-l))(e(\beta(l+t))-1)}{l+t}\overline{\chi}(k-l).\end{split}$
@@
$\frac{2\pi ie(-\alpha\left\{q\theta\right\})}{\tau(\chi)}\sum_{\alpha q\leq n\leq\alpha q+\beta q}\chi(n)e(n\theta)=\sum_{l\in\mathbb{Z}}\frac{e(\alpha l)(e(\beta(l+t))-1)}{l+t}\overline{\chi}(k-l).$
@@
We will first prove that we can truncate the infinite sum over $l$ when we vary either the character or the shift. We choose some slowly growing function $L(q)$ and for each $k$ and $\chi$, we define the function
@@
Let $L(q)$ be some function such that $L(q)\rightarrow\infty$ and $L(q)\ll q^{\varepsilon}$ for all $\varepsilon>0$. For fixed real numbers $\alpha$ and $\beta$ and for any bounded Lipschitz function $h:\mathscr{C}[0,1]\rightarrow\mathbb{R}$, and any sequence of integers $k_{q}$, the expected value $\operatorname{\mathbb{E}}h(F_{k_{q},q,\alpha,\beta})$ approaches $\operatorname{\mathbb{E}}h(\tilde{F}_{k_{q},q,\alpha,\beta})$ as $q$ tends to infinity over the primes, in the sense that
$\left|\operatorname{\mathbb{E}}h(F_{k_{q},q,\alpha,\beta})-\operatorname{\mathbb{E}}h(\tilde{F}_{k_{q},q,\alpha,\beta})\right|\rightarrow 0.$
That is, $F_{k_{q},q,\alpha,\beta}$ converges weakly to some $F$ (as $q$ tends to infinity over the primes, or over some subsequence of the primes) if and only if $\tilde{F}_{k_{q},q,\alpha,\beta}$ converges weakly to $F$.
Similarly, the expected value $\operatorname{\mathbb{E}}h(G_{q,\alpha,\beta})$ approaches $\operatorname{\mathbb{E}}h(\tilde{G}_{q,\alpha,\beta})$.
@@
We will abbreviate $F_{k,\chi}:=F_{k,\chi,\alpha,\beta}$. We start simply with the triangle inequality.
@@ 4. Tightness and continuity
The last sum is absolutely convergent, hence continuous for any realization of $\mathbb{W}(l)$. So the only problems with convergence may arise from the first two sums. But moment bounds easily show that they converge with probability 1. More generally, we will shortly compute moment bounds for all finite dimensional distributions of $F_{k,\alpha,\beta}$ which imply almost-sure convergence for every $t$. ∎
@@ 5.1. Varying character $\chi{\text{\rm\ (mod~$q$)}}$
$F(t)=\frac{e(\alpha t)}{2\pi i}\sum_{l\in\mathbb{Z}}\frac{e(\beta(l+t))-1}{l+t}\mathbb{X}(n).$
@@
with $n_{l},m_{l}<L(q)$ are the “trivial” ones (those where $m_{1},\ldots,m_{r}$ is a permutation of $n_{1},\ldots,n_{r}$). In general, this may not be the case. For each $n_{1},\ldots,n_{r}$ and $m_{1},\ldots,m_{r}$ we consider the polynomial
@@
If $\mathbf{n}$ is not a permutation of $\mathbf{m}$, this is not the zero polynomial. The coefficients are bounded by some power of $L(q)$, and it is exactly the content of Lemma 7 that there exist choices of $L(q)$ such that: in case (1), there are only finitely many primes such that $h(k_{q})\equiv 0\pmod{q}$ for any non-zero polynomial with small coefficients; in each of cases (2) and (3), there is a full density subset of the primes $\mathcal{P}$ such that $h(k_{q})\neq 0\pmod{q}$ for any non-zero polynomial with small coefficients when $q\in\mathcal{P}$ is large enough. ∎
@@ 5.2. Quadratic characters
$\frac{q^{1/2}}{\tau(\chi)}F_{k,\chi_{q},\alpha,\beta}(t)=\frac{e(-\alpha k)}{\tau(\chi)}\sum_{\alpha q<n<(\alpha+\beta)q}\chi_{q}(n)e(n(k+t)/q),$
@@
$\frac{q^{1/2}}{\tau(\chi)}\tilde{F}_{k,\chi_{q},\alpha,\beta}(t)=\frac{e(\alpha t)}{2\pi i}\sum_{\left|l\right|<L(q)}\frac{e(\alpha l)(e(\beta(l+t))-1)}{l+t}\chi_{q}(k-l).$
@@
We begin with the same notation and combinatorics as in the proof of Proposition 5, recalling Equation (7):
$\prod_{j=1}^{J}\frac{q^{1/2}}{\tau(\chi)}\tilde{F}_{k,\chi_{q},\alpha,\beta}(t_{j})^{r_{j}}\overline{\tilde{F}_{k,\chi_{q},\alpha,\beta}(t_{j})}^{s_{j}}\\ =e(\alpha T)\frac{1}{(2\pi)^{r+s}}\sum_{\mathbf{m}_{1},\dots,\mathbf{m}_{J}}\prod_{j=1}^{J}\prod_{\ell=1}^{r_{j}}\chi_{q}(k+m_{j,\ell})\prod_{\ell=r_{j}+1}^{s_{j}}\chi_{q}(k+m_{j,\ell})\prod_{j=1}^{J}a(\mathbf{m}_{j};t_{j}).$
We take the sum over $k$ to compute the expectation.
@@ 6. Prime removal lemmas
    $\left\{p\leq X:f(k_{p})\equiv 0{\text{\rm\ (mod~$p$)}}\right\}\ll d\log\max\left|e_{j}\right|\psi(X).$
@@
We now notice that $qak\equiv-ar{\text{\rm\ (mod~$kq+r$)}}$, removing the dependence on $k$:
@@
Now we notice that $\sum_{j=0}^{d}e_{j}(-ar+qc+q\varepsilon)^{j}q^{d-j}$ is just a number. For a given $X$ it depends only on $r$ and $\varepsilon$; it is bounded by $O(\max\left|e_{j}\right|(q+X/Q)^{2d})$ because $a,r,$ and $c$ are <q<q and $\varepsilon\ll X/qQ$; it is nonzero because it is $\equiv-ar{\text{\rm\ (mod~$q$)}}$ and $a$ and $r$ are coprime to $q$. So it is divisible by at most
@@
We finish by making the optimal choice $Q=X^{1/2}$. ∎
@@
    For any sequence $k_{q}\to\infty$ such that $k_{q}=q^{o(1)}$, given any integer $d>0$ and for all large enough (in terms of $d$) primes $q$, any polynomial $f$ of degree $d$ with coefficients bounded by $L(q)^{d+1}$, satisfies $f(k_{q})\not\equiv 0\pmod{q}$.
@@
    For any sequence $k_{q}\to\infty$ such that $k_{q}<\psi(q)$ for an increasing function $\psi$ with $\psi(X)=o(\pi(X))$, there exists a subset of the primes $\mathcal{P}$ such that
    • $(\mathcal{P}_{\theta}\cap[1,X])=\pi(X)(1+o(1))$ for all $X$,
    • for every positive integer $d$ there are only finitely many $q\in\mathcal{P}$ (in terms of $d$) such that there exists a polynomial $f$ of degree $d$ with coefficients bounded by $L(q)^{d+1}$ with $f(k_{q})\equiv 0{\text{\rm\ (mod~$q$)}}$.
@@
    • $(\mathcal{P}_{\theta}\cap[1,X])>\pi(X)-O(X^{1/2+\varepsilon})$ for all $X$,
@@
where $f:[\alpha,\beta]\to\mathbb{R}$ is a smooth function, when $\varepsilon$ is small. Note that if this quantity goes to 0 as $\varepsilon\to 0$, this guarantees the existence of $\int_{\alpha}^{\beta}\log\lvert f(t)\rvert\mathrm{d}t.$ In this section, we are going to prove the following key result.
@@
By Lemma 8 and the condition $\Delta f(t)\geq 1$ for all $t\in[\alpha,\beta]$, $f$ has at most $T$ zeros, for some finite number $T$ (independent of $f$). For each zero $x$ of $f$, let $(x)$ denote the interval containing $x$ such that $\lvert f(x)\rvert<\varepsilon$.
@@
(9) $\beta(t)=\begin{cases}e^{-1/t},&t>0,\\[5.69054pt] 0,&t\leq 0,\end{cases}\qquad\rho(t)=\frac{\beta(t)}{\beta(t)+\beta(1-t)},$
@@ 8.1. Proof of Theorem 3
The main difficulty in proving Theorem 3 is that the functional on $\mathscr{C}[0,1]$ defined by $\ell(f)=\int_{0}^{1}\log(|f(t)|)\text{d}t$ is not continuous, so our previous distributional result does not directly apply. We instead consider
$\ell(f)=\ell_{\varepsilon}(f)+\int_{0}^{1}\log(|f(t)|)(1-w_{\varepsilon}(\lvert f(t)\rvert))\mathrm{d}t,$
@@
$\ell_{\varepsilon}(f)=\int_{0}^{1}\log(|f(t)|)w_{\varepsilon}(\lvert f(t)\rvert)\text{d}t$
with $w_{\varepsilon}(t)=\rho\left(\frac{t}{\varepsilon}-1\right)$ (recall the function $\rho$ in (9)) a smoothed minorant of $\mathds{1}_{t\geq\varepsilon}$, is a continuous functional on $\mathscr{C}[0,1]$. Roughly speaking, Theorem 3 will follow from applying Theorem 2 to $\ell_{\varepsilon}(f)$, and Proposition 8 to $\ell(f)-\ell_{\varepsilon}(f)$ using the fact that $1-w_{\varepsilon}(t)$ majorizes $\mathds{1}_{t<\varepsilon}$, where $f=F_{k,\chi_{q},\alpha,\beta}$ with $\alpha=.2$ and $\beta=1.1$.
@@
For $\alpha=.2$ and $\beta=1.1$, for all large enough primes $q$ and all $k\in\{0,1,\ldots,q-1\}$ we have
@@
In any case, we find that we can always cover $[0,1]$ by 5 intervals such that on each interval, either $\left|\mathrm{Re}(4\pi^{2}\beta^{2}A^{\prime}(t)+A^{\prime\prime\prime}(t))\right|>1$ or $\left|\mathrm{Im}(4\pi^{2}\beta^{2}A^{\prime}(t)+A^{\prime\prime\prime}(t))\right|>1$ across the entire interval.
@@
Applying Theorem 2 to the continuous functional $\ell_{\varepsilon}$, and using Lemma 10 together with the fact that $1-w_{\varepsilon}(t)$ majorizes $\mathds{1}_{t<\varepsilon}$ (note that $\log\lvert t\rvert$ is negative for small enough $t$), we have
$\int_{0}^{1}\log\left|\frac{1}{(1.1q)^{1/2}}S(\chi_{q},.2,1.1,\theta)\right|\mathrm{d}\theta\longrightarrow\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}G_{.2,1.1}(t)\right|w_{\varepsilon}(\lvert G_{.2,1.1}(t)\rvert)+O(\varepsilon^{1/6}).$
@@
$\int_{0}^{1}\log\left|\frac{1}{(1.1q)^{1/2}}S(\chi_{q},.2,1.1,\theta)\right|\mathrm{d}\theta\longrightarrow\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}G_{.2,1.1}(t)\right|\mathrm{d}t.$
The exponential of the latter expectation appears to be equal to $954\ldots.$

### v2

@@ 2.1. Proofs of Theorems 1 and 2
for all (sufficiently large) $k$ and for all $s,t\in[0,1]$, and if the finite dimensional distributions of $L_{k}(t)$ tend to some limits, then $L_{k}(t)$ converges weakly to some random process $L$.
Instead of working directly with the random processes $F_{k,q,\alpha,\beta}$ and $G_{q,\alpha,\beta}$, we define truncated versions $\tilde{F}_{k,q,\alpha,\beta}$ and $\tilde{G}_{q,\alpha,\beta}$ and prove in Proposition 4 that a sequence of these truncated random processes converge if and only if the untruncated processes converge. This is a technical convenience, but it also somewhat necessary; the moments of the untruncated random processes will not converge without removal of the trivial character and moreover, the size of the truncation parameter will play a role in our counting arguments handling various ranges of $k_{q}.$
@@
Finally, to apply Proposition 2, we prove that all mixed moments of all of finite dimentional distributions of $\tilde{F}_{k,q,\alpha,\beta}$ tend to those of $F_{k,\alpha,\beta}$ when $k$ is fixed (Proposition 5) or tend to those of $F_{\alpha,\beta}$ when $k$ tends to infinity, after possibly removing some primes (Proposition 6), in each of the cases (2), (3), and (4) of Theorem 1. These propositions also include bounds for the moments which show that they determine the distribution (Carleman’s condition), yielding a proof of Theorem 1. Similarly, we deal with the quadratic case in Proposition 7, completing the proof of Theorem 2.
@@ 3. The random processes and truncation
$\tilde{f}(n)=\frac{1}{2}\lim_{\varepsilon\rightarrow 0}\big(f(n+\varepsilon)+f(n-\varepsilon)\big)$
and $\tau(\chi)$ denotes the Gauss sum of $\chi$.
@@
Equipped with these lemmas we are in a position to state and prove our general formula for incomplete shifted mixed character sums.
@@
For a primitive Dirichlet character $\chi$ mod $q$ and for $\alpha,\beta\in\mathbb{R}$, we have
@@
In particular, for any $\alpha$ and $\beta$ such that $\alpha q$ and $(\alpha+\beta)q$ are not integers, we have
@@
$\sideset{}{{}^{\prime}}{\sum}_{\alpha q\leq n\leq\alpha q+\beta q}\chi(n)e(n\theta)=e(\alpha q\theta)\frac{\tau(\chi)}{2\pi i}\sum_{m\in\mathbb{Z}}\frac{e(-\alpha m)(e(\beta(\theta q-m))-1)}{\theta q-m}\overline{\chi}(m),$
where $\sum^{\prime}$ indicates that if $\alpha q$ or $(\alpha+\beta)q$ is an integer, then it is counted with a $\frac{1}{2}$-weight. Now write $\theta=(k+t)/q$, with $k=\left\lfloor q\theta\right\rfloor$ and $t=\left\{\theta\right\}$. Then
$\begin{split}e(-\alpha q\theta)\sideset{}{{}^{\prime}}{\sum}_{\alpha q\leq n\leq\alpha q+\beta q}\chi(n)e(n\theta)=&\frac{\tau(\chi)}{2\pi i}\sum_{m\in\mathbb{Z}}\frac{e(-\alpha m)(e(\beta(k+t-m))-1)}{k+t-m}\overline{\chi}(m)\\ =&\frac{\tau(\chi)}{2\pi i}\sum_{l\in\mathbb{Z}}\frac{e(-\alpha(k-l))(e(\beta(l+t))-1)}{l+t}\overline{\chi}(k-l).\end{split}$
@@
$\frac{2\pi ie(-\alpha\left\{q\theta\right\})}{\tau(\chi)}\sideset{}{{}^{\prime}}{\sum}_{\alpha q\leq n\leq\alpha q+\beta q}\chi(n)e(n\theta)=\sum_{l\in\mathbb{Z}}\frac{e(\alpha l)(e(\beta(l+t))-1)}{l+t}\overline{\chi}(k-l).$
@@
We will first prove that we can truncate the infinite sum over $l$ when we vary either the character $\chi$ or the shift $k$. We choose some slowly growing function $L(q)$ and for each $k$ and $\chi$, we define the function
@@
Let $L(q)$ be some function such that $L(q)\rightarrow\infty$ and $L(q)\ll q^{\varepsilon}$ for all $\varepsilon>0$. For fixed real numbers $\alpha$ and $\beta$, any bounded Lipschitz function $h:\mathscr{C}[0,1]\rightarrow\mathbb{R}$, and any sequence of integers $k_{q}$, as $q$ tends to infinity over the primes we have
$\left|\operatorname{\mathbb{E}}h(F_{k_{q},q,\alpha,\beta})-\operatorname{\mathbb{E}}h(\tilde{F}_{k_{q},q,\alpha,\beta})\right|\rightarrow 0\text{ and }\left|\operatorname{\mathbb{E}}h(G_{q,\alpha,\beta})-\operatorname{\mathbb{E}}h(\tilde{G}_{q,\alpha,\beta})\right|\rightarrow 0.$
That is, $F_{k_{q},q,\alpha,\beta}$ converges weakly to some $F$ (as $q$ tends to infinity over the primes, or over some subsequence of the primes) if and only if $\tilde{F}_{k_{q},q,\alpha,\beta}$ converges weakly to $F$, and similarly for $G_{q,\alpha,\beta}$ and $\tilde{G}_{q,\alpha,\beta}$.
@@
We will abbreviate $F_{k,\chi}:=F_{k,\chi,\alpha,\beta}$. By the triangle inequality
@@ 4. Tightness and continuity
The last sum is absolutely convergent, hence continuous for any realization of $\mathbb{W}(l)$. So the only problems with convergence may arise from the first two sums, which are independent of $t$. But moment bounds easily show that they converge with probability 1. More generally, we will shortly compute moment bounds for all finite dimensional distributions of $F_{k,\alpha,\beta}$ which imply almost-sure convergence for every $t$. ∎
@@ 5.1. Varying character $\chi{\text{\rm\ (mod~$q$)}}$
$F(t)=\frac{e(\alpha t)}{2\pi i}\sum_{l\in\mathbb{Z}}\frac{e(\beta(l+t))-1}{l+t}\mathbb{X}(l).$
@@
with $n_{l},m_{l}<L(q)$ are those where $m_{1},\ldots,m_{r}$ is a permutation of $n_{1},\ldots,n_{r}$. In general, this may not be the case. For each $n_{1},\ldots,n_{r}$ and $m_{1},\ldots,m_{r}$ we consider the polynomial
@@
If $\mathbf{n}$ is not a permutation of $\mathbf{m}$, this is not the zero polynomial. The coefficients are bounded by some power of $L(q)$, and it is exactly the content of Lemma 7 (below in Section 6) that there exist choices of $L(q)$ such that: in case (1), there are only finitely many primes such that $h(k_{q})\equiv 0\pmod{q}$ for any non-zero polynomial with small coefficients; in each of cases (2) and (3), there is a full density subset of the primes $\mathcal{P}$ such that $h(k_{q})\not\equiv 0\pmod{q}$ for any non-zero polynomial with small coefficients when $q\in\mathcal{P}$ is large enough. ∎
@@ 5.2. Quadratic characters
$\frac{q^{1/2}}{\tau(\chi_{q})}F_{k,\chi_{q},\alpha,\beta}(t)=\frac{e(-\alpha k)}{\tau(\chi_{q})}\sum_{\alpha q<n\leq(\alpha+\beta)q}\chi_{q}(n)e(n(k+t)/q),$
@@
$\frac{q^{1/2}}{\tau(\chi_{q})}\tilde{F}_{k,\chi_{q},\alpha,\beta}(t)=\frac{e(\alpha t)}{2\pi i}\sum_{\left|l\right|<L(q)}\frac{e(\alpha l)(e(\beta(l+t))-1)}{l+t}\chi_{q}(k-l).$
@@
We begin with the same notation and combinatorics as in the proof of Proposition 5, recalling (7):
$\prod_{j=1}^{J}\frac{q^{1/2}}{\tau(\chi_{q})}\tilde{F}_{k,\chi_{q},\alpha,\beta}(t_{j})^{r_{j}}\overline{\tilde{F}_{k,\chi_{q},\alpha,\beta}(t_{j})}^{s_{j}}\\ =e(\alpha T)\frac{1}{(2\pi)^{r+s}}\sum_{\mathbf{m}_{1},\dots,\mathbf{m}_{J}}\prod_{j=1}^{J}\prod_{\ell=1}^{r_{j}}\chi_{q}(k+m_{j,\ell})\prod_{\ell=r_{j}+1}^{s_{j}}\chi_{q}(k+m_{j,\ell})\prod_{j=1}^{J}a(\mathbf{m}_{j};t_{j}).$
We take the sum over $k$ to compute the expectation
@@ 6. Prime removal lemmas
    $\left\{p\leq X:f(k_{p})\equiv 0{\text{\rm\ (mod~$p$)}}\right\}\ll d\psi(X)\log\max\left|e_{j}\right|.$
@@
We now notice that $qak\equiv-ar{\text{\rm\ (mod~$kq+r$)}}$, removing the dependence on $k$, that is
@@
Next notice that $\sum_{j=0}^{d}e_{j}(-ar+qc+q\varepsilon)^{j}q^{d-j}$ is just a number, which for a given $X$ depends only on $r$ and $\varepsilon$. It is bounded by $O(\max\left|e_{j}\right|(q+X/Q)^{2d})$ because $a,r,$ and $c$ are <q<q and $\varepsilon\ll X/qQ$, and it is nonzero because it is $\equiv-ar{\text{\rm\ (mod~$q$)}}$ and $a$ and $r$ are coprime to $q$. So it is divisible by at most
@@
We finish by making the essentially optimal choice $Q=X^{1/2}$, giving the claimed bound. ∎
@@
    For any sequence $k_{q}\to\infty$ such that $k_{q}=q^{o(1)}$, given any positive integer $d$ and for all large enough (in terms of $d$) primes $q$, any polynomial $f$ of degree $d$ with coefficients bounded by $L(q)^{d+1}$, satisfies $f(k_{q})\not\equiv 0\pmod{q}$.
@@
    For any sequence $k_{q}\to\infty$ such that $k_{q}<\psi(q)$ for an increasing function $\psi$ with $\psi(X)=o(\pi(X))$, there exists a subset of the primes $\mathcal{P}$ such that:
    • $(\mathcal{P}_{\theta}\cap[1,X])=\pi(X)(1-o(1))$ for all $X$,
    • for every positive integer $d$ there are only finitely many (in terms of $d$) $q\in\mathcal{P}$ such that there exists a polynomial $f$ of degree $d$ with coefficients bounded by $L(q)^{d+1}$ with $f(k_{q})\equiv 0{\text{\rm\ (mod~$q$)}}$.
@@
    • $(\mathcal{P}_{\theta}\cap[1,X])>\pi(X)-O_{\varepsilon}(X^{1/2+\varepsilon})$ for all $X$,
@@
uniformly in smooth functions $f:[\alpha,\beta]\to\mathbb{R}$, when $\varepsilon$ is small. In this section, we are going to prove the following key result.
@@
[…]
@@
(9) $\beta(t)=\begin{cases}e^{-1/t}&t>0,\\[5.69054pt] 0&t\leq 0,\end{cases}\qquad\rho(t)=\frac{\beta(t)}{\beta(t)+\beta(1-t)},$
@@ 8.1. Proof of Theorem 3
The main difficulty in proving Theorem 3 is that the functional on $\mathscr{C}[0,1]$ defined by $\ell(f)=\int_{0}^{1}\log\left|f(t)\right|\mathrm{d}t$ is not continuous, so our previous distributional result does not directly apply. We instead consider
$\ell(f)=\ell_{\varepsilon}(f)+\int_{0}^{1}\log\left|f(t)\right|(1-w_{\varepsilon}(\lvert f(t)\rvert))\mathrm{d}t,$
@@
$\ell_{\varepsilon}(f)=\int_{0}^{1}\log\left|f(t)\right|w_{\varepsilon}(\lvert f(t)\rvert)\mathrm{d}t$
with $w_{\varepsilon}(t)=\rho\left(\frac{t}{\varepsilon}-1\right)$ (recall the function $\rho$ in (9)) a smoothed minorant of $\mathds{1}_{t\geq\varepsilon}$, is a continuous functional on $\mathscr{C}[0,1]$. Roughly speaking, Theorem 3 will follow from applying Theorem 2 to a bounded version of $\ell_{\varepsilon}(f)$, and Proposition 8 to $\ell(f)-\ell_{\varepsilon}(f)$ using the fact that $1-w_{\varepsilon}(t)$ majorizes $\mathds{1}_{t<\varepsilon}$, where $f=F_{k,\chi_{q},\alpha,\beta}$ with $\alpha=0.2$ and $\beta=1.1$. Indeed, considering a bounded version will be enough, for we will show that $\ell_{\varepsilon}(F_{k,\chi_{q},\alpha,\beta})$ takes large values with small probability; see Lemma 11 below.
@@
For $\alpha=0.2$ and $\beta=1.1$, for all large enough primes $q$ and all $k\in\{0,1,\ldots,q-1\}$ we have
@@
In any case, we find that we can always cover $[0,1]$ by five intervals such that on each interval, either $\left|\mathrm{Re}(4\pi^{2}\beta^{2}A^{\prime}(t)+A^{\prime\prime\prime}(t))\right|>1$ or $\left|\mathrm{Im}(4\pi^{2}\beta^{2}A^{\prime}(t)+A^{\prime\prime\prime}(t))\right|>1$ across the entire interval.
@@
Now we have
(10) $\int_{0}^{1}\log\left|\frac{1}{(1.1q)^{1/2}}S(\chi_{q},0.2,1.1,\theta)\right|\mathrm{d}\theta=\\ \mathbb{E}\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}G_{q,0.2,1.1}(t)\ \right\rvert(w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert)+(1-w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert))\mathrm{d}t.$
Lemma 11.
For any small enough $\varepsilon>0$, we have
[…]
Proof.
Note that for every realization of $G_{q,0.2,1.1}$, we have
$\left(\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}G_{q,0.2,1.1}(t)\right\rvert w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert)\mathrm{d}t\right)^{2}\ll_{\varepsilon}\int_{0}^{1}\left\lvert G_{q,0.2,1.1}(t)\right\rvert^{2}\mathrm{d}t$
since Mahler measure is bounded above by the $L_{2}$ norm, whereas for the lower bound it suffices to take the implied constant to be $(\log(1/\varepsilon)/\varepsilon)^{2}$ since the integrand is supported on $\lvert G_{q,0.2,1.1}(t)\rvert\gg\varepsilon$. Taking expectation of both sides of the above inequality, considering that $\varepsilon>0$ is fixed, we obtain
$\mathbb{E}\left(\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}G_{q,0.2,1.1}(t)\ \right\rvert w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert)\mathrm{d}t\right)^{2}\ll 1,$
since
$\mathbb{E}\int_{0}^{1}\left\lvert G_{q,0.2,1.1}(t)\right\rvert^{2}\mathrm{d}t$ $=\frac{1}{q}\int_{0}^{1}\sum_{\alpha q<n_{1},n_{2}\leq(\alpha+\beta)q}\left(\frac{n_{1}n_{2}}{q}\right)\mathbb{E}e\left(\frac{k+t}{q}(n_{1}-n_{2})\right)\mathrm{d}t$
$=\frac{1}{q}\int_{0}^{1}\sum_{\begin{subarray}{c}\alpha q<n_{1},n_{2}\leq(\alpha+\beta)q\\ q\mid n_{1}-n_{2}\end{subarray}}e\left(\frac{t(n_{1}-n_{2})}{q}\right)\mathrm{d}t$
$\ll 1,$
since only the terms $n_{1}=n_{2}$ give a non-zero contribution, and there are $\ll\beta q$ such terms. Markov’s inequality gives
$\mathbb{P}\left(\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}G_{q,0.2,1.1}(t)\right\rvert w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert)\mathrm{d}t\geq M\right)\ll\frac{1}{M^{2}}.$
Therefore for any large $M>0$ we have
$\mathbb{E}\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}G_{q,0.2,1.1}(t)\right\rvert w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert)\mathrm{d}t\\ =\mathbb{E}\min\left\{\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}G_{q,0.2,1.1}(t)\right\rvert w_{\varepsilon}(\lvert G_{q,0.2,1.1}(t)\rvert)\mathrm{d}t,M\right\}+O\left(\frac{1}{M}\right).$
Applying Theorem 2 to the bounded continuous functional
$f\mapsto\min\left\{\int_{0}^{1}\log\left\lvert\frac{1}{1.1^{1/2}}f(t)\right\rvert w_{\varepsilon}(\lvert f(t)\rvert)\mathrm{d}t,M\right\}$
and letting $M\to\infty$ (with fixed $\varepsilon>0$) together with an application of the monotone convergence theorem yields the desired result. ∎
Next, returning to (10), using Lemmas 11 and 10 together with the fact that $1-w_{\varepsilon}(t)$ majorizes $\mathds{1}_{t<\varepsilon}$ (note that $\log\lvert t\rvert$ is negative for small enough $t$), we have
$\int_{0}^{1}\log\left|\frac{1}{(1.1q)^{1/2}}S(\chi_{q},0.2,1.1,\theta)\right|\mathrm{d}\theta\longrightarrow\\ \mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}G_{0.2,1.1}(t)\right|w_{\varepsilon}(\lvert G_{0.2,1.1}(t)\rvert)\mathrm{d}t+O(\varepsilon^{1/6}).$
@@
$\int_{0}^{1}\log\left|\frac{1}{(1.1q)^{1/2}}S(\chi_{q},0.2,1.1,\theta)\right|\mathrm{d}\theta\longrightarrow\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}G_{0.2,1.1}(t)\right|\mathrm{d}t.$
The exponential of the latter expectation may be computed to be equal to $0.954\ldots$.